A boy is 2 years more than twice as old as his younger brother.The two boys together are 17 years older than their sister who is 7 years younger than the older boy.How old is the sister?
Answer by JulietG(1812)

Let A be the older boy and B be the younger.
A = 2B + 2 [A boy is 2 years more than twice as old as his younger brother.]
A + B = S + 17 [The two boys together are 17 years older than their sister]
A = S + 7 [who is 7 years younger than the older boy]
.
There are several ways to do this. Here is one:
A + B = S + 17
(S + 7) + B = S + 17
S + B = S + 10
B = 10
The younger boy is 10. Knowing this, we can find the others.
A = 2B + 2
A = 2(10) + 2
A = 20 + 2
A = 22
The older boy is 22
The sister is A = S + 7
22 = S + 7
15 = Sister